Blood flows in an aorta of radius 1.0 cm at 30cm/s. What is the volume flow rate?


1
Expert's answer
2022-02-14T14:59:24-0500

"V=\pi r^2h=\pi *1^2*30=30\pi\approx 94.25 \frac{cm^3}{s}" .